OpenStudy (anonymous):

(-9 / 3) / ((-6)^2 / 12) @JoStudy

OpenStudy (anonymous):

OK So is that \[\huge \frac{ -9 }{ 3 }/\frac{ -6^{2} }{ 12 }\]

OpenStudy (anonymous):

yepp

OpenStudy (anonymous):

OK So rememebr PEDMAS what do you solve first in this equation?

OpenStudy (anonymous):

-6^2/12

OpenStudy (anonymous):

Ah close but out of that equation which do you solve first?

OpenStudy (anonymous):

-9/3

OpenStudy (anonymous):

Think about it pEdmas

OpenStudy (anonymous):

\(\Huge \frac{ -9 }{ 3 }/\frac{\color{red}{ -6^{2}} }{ 12 }\)

OpenStudy (anonymous):

6^2

OpenStudy (anonymous):

-6^2

OpenStudy (anonymous):

-9/3 / -36/12

OpenStudy (anonymous):

OK\ \[\huge\frac{ -9 }{3}\ \frac{-36}{12}\]

OpenStudy (anonymous):

Do you know how to divide fractions?

OpenStudy (anonymous):

did u switch it up??

OpenStudy (anonymous):

Yup :) you invert one of the fractions then multiply

OpenStudy (anonymous):

so it would be -9/3 / 12/-6 ??

OpenStudy (anonymous):

-9/3 / 12/-36

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

\[\huge\frac{ -9 }{ 3 }*\frac{ 12}{ -36 }\]

OpenStudy (anonymous):

I got 1

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

u make it so much easier

OpenStudy (anonymous):

Thanks :) I try to make it understandable :)